# Multi-Level Multi-Fidelity Methods for Path Integral and Safe Control

### v1

@@ 3 Problem Formulation
where $N$ is the number of trajectories and $\mathds{1}(\cdot)$ is the indicator function.
@@
where the sum $\sum$ is taken over $N$ simulated trajectories, and $x^{n},u^{n},w^{n}$ are realizations of the state, control, and noise for the $n$-th trajectory. Intuitively, each trajectory is assigned a weight $\exp[-\int_{t}^{T}g(x_{\tau}^{n},u_{\tau}^{n})\,d\tau-\Psi(x_{T}^{n})]$, which encodes both the running cost $g(\cdot)$ along the path and the terminal cost $\Psi(\cdot)$. Trajectories that yield lower cost receive higher weight, while high-cost trajectories are exponentially suppressed. The numerator in (8) collects weighted fluctuations of the noise increments $\int_{t}^{s}dw_{\tau}^{n}$, and the denominator normalizes these contributions to ensure a proper probability-weighted average. We refer readers to [11] for more detailed discussions on path integral control.
@@ 4.1 Multi-Level Monte Carlo
Denote $P$ the value of interest to be estimated, and $P_{1},P_{2},\cdots,P_{L}$ are estimates from level $1$ to $L$ with increasing accuracy but also increasing cost. The MLMC estimation is given by
@@ 4.3 Multi-level Multi-fidelity Monte Carlo
The parameters $a_{L}=1$ will ensure unbiased MLMF estimator under certain conditions, as shown later in Theorem 5.3. Different configurations of $a_{l}$ for $l=1,\cdots L-1$ will result in different variances at each level, which affect the sample complexity of the estimator. Optimal $a_{l}$ are derived in Theorem 5.4 and the corresponding estimation variance is derived in Theorem 5.5.
@@ 7.1 Safety Probability Estimation
where $\mathbf{x}=[x_{1},x_{2},\cdots,x_{5}]^{\top}\in\mathbb{R}^{5}$ is the state, $W_{t}$ is the 5-dimensional standard Wiener process with $W_{0}=\boldsymbol{0}$, $K=[K_{1},K_{2},K_{3},K_{4},K_{5}]=[1,0.7,0.6,0.5,0.4]$ is the system parameter and $\sigma=[\sigma_{1},\sigma_{2},\sigma_{3},\sigma_{4},\sigma_{5}]=[1,0.7,0.6,0.5,0.4]$ is the noise magnitude. The safe set is given by $\mathcal{C}=\{\mathbf{x}:x_{i}\geq 0,\forall i=1,2,\cdots,5\}$. The computation cost for sampling each dimension $i\in\{1,2,3,4,5\}$ is assumed to be $C_{1}=1,\;C_{2}=5,\;C_{3}=10,\;C_{4}=100,\;C_{5}=1000$.
We set the initial state to be $\mathbf{x}_{0}=[1,1,\cdots,1]^{\top}$, and estimate the safety probability $F(\mathbf{x}_{0})$ defined in (3) with horizon $T=1$ using the proposed MLMF method with $a_{l}\equiv 1$, $\Delta t_{l}=0.05$ for $l=1,2,3$ and $\Delta t_{l}=0.1$ for $l=4,5$. We obtain each level of estimate by simulating the first $l$ dimensions of the system (i.e., $\xi^{(l)}=[x_{1},\cdots,x_{l}]^{\top}$) and calculating the safety probability with regard to safe set $\mathcal{C}_{l}=\{\xi^{(l)}:x_{i}\geq 0,\forall i=1,\cdots,l\}$. We compare with (i) standard MC where we directly estimate the safety probability with the 5-dimensional dynamics, (ii) subset simulation where we multiply conditional probabilities of failure for transitional safe set $\mathcal{C}_{p}=\{\mathbf{x}:x_{i}\geq p,\forall i=1,\cdots,5\}$ with $p=\{0.8,0.5,0.3,0.2,0\}$, as well as (iii) MLMC (layering in time only) and (iv) MFMC (layering in state space only). Fig. 2 shows the absolute error of the estimate with regard to computation cost for all methods. The ground truth value is obtained by running standard MC with $10^{8}$ samples. We can see that with more samples, the absolute errors decrease for all methods. Layering in time (MLMC) or in state (MFMC) alone results in error reduction compared to standard MC and subset simulation. By leveraging both, the proposed MLMF outperforms all other methods.

### v2

@@ 3 Problem Formulation
where NN is the number of trajectories and $\mathds{1}(\cdot)$ is the indicator function.
@@
where the sum $\sum$ is taken over NN simulated trajectories, and $x^{n},u^{n},w^{n}$ are realizations of the state, control, and noise for the $n$-th trajectory. Intuitively, each trajectory is assigned a weight $\exp[-\int_{t}^{T}g(x_{\tau}^{n},u_{\tau}^{n})\,d\tau-\Psi(x_{T}^{n})]$, which encodes both the running cost $g(\cdot)$ along the path and the terminal cost $\Psi(\cdot)$. Trajectories that yield lower cost receive higher weight, while high-cost trajectories are exponentially suppressed. The numerator in (8) collects weighted fluctuations of the noise increments $\int_{t}^{s}dw_{\tau}^{n}$, and the denominator normalizes these contributions to ensure a proper probability-weighted average. We refer readers to [11] for more detailed discussions on path integral control.
@@ 4.1 Multi-Level Monte Carlo
Denote $P$ the value of interest to be estimated, and $P_{1},P_{2},\cdots,P_{L}$ are estimates from level 11 to $L$ with increasing accuracy but also increasing cost. The MLMC estimation is given by
@@ 4.3 Multi-level Multi-fidelity Monte Carlo
The parameters $a_{L}=1$ will ensure unbiased MLMF estimator under certain conditions, as shown later in Theorem 5.3. Different configurations of $a_{l}$ for $l=1,\cdots L-1$ will result in different variances at each level, which affect the sample complexity of the estimator. Optimal $a_{l}$ are derived in Theorem 5.4 and the corresponding estimation variance is derived in Theorem $5.$
@@ 7.1 Safety Probability Estimation
where $\mathbf{x}=[x_{1},x_{2},\cdots,x_{5}]^{\top}\in\mathbb{R}^{5}$ is the state, $W_{t}$ is the 5-dimensional standard Wiener process with $W_{0}=\boldsymbol{0}$, $K=[K_{1},K_{2},K_{3},K_{4},K_{5}]=[1,0.7,0.6,0.5,0.4]$ is the system parameter and $\sigma=[\sigma_{1},\sigma_{2},\sigma_{3},\sigma_{4},\sigma_{5}]=[1,0.7,0.6,0.5,0.4]$ is the noise magnitude. The safe set is given by $\mathcal{C}=\{\mathbf{x}:x_{i}\geq 0,\forall i=1,2,\cdots,5\}$. The computation cost for sampling each dimension $i\in\{1,2,3,4,5\}$ is assumed to be $C_{1}=1,\;C_{2}=10,\;C_{3}=20,\;C_{4}=50,\;C_{5}=500$.
We set the initial state to be $\mathbf{x}_{0}=[1,1,\cdots,1]^{\top}$, and estimate the safety probability $F(\mathbf{x}_{0})$ defined in (3) with horizon $T=1$ using the proposed MLMF method with $a_{l}\equiv 1$, $\Delta t_{l}=0.1$ for $l=1,2,3$ and $\Delta t_{l}=0.05$ for $l=4,5$. We obtain each level of estimate by simulating the first $l$ dimensions of the system (i.e., $\xi^{(l)}=[x_{1},\cdots,x_{l}]^{\top}$) and calculating the safety probability with regard to safe set $\mathcal{C}_{l}=\{\xi^{(l)}:x_{i}\geq 0,\forall i=1,\cdots,l\}$. We compare with (i) standard MC where we directly estimate the safety probability with the 5-dimensional dynamics, (ii) subset simulation where we multiply conditional probabilities of failure for transitional safe set $\mathcal{C}_{p}=\{\mathbf{x}:x_{i}\geq p,\forall i=1,\cdots,5\}$ with $p=\{0.8,0.5,0.3,0.2,0\}$, as well as (iii) MLMC (layering in time only) and (iv) MFMC (layering in state space only). Fig. 2 shows the absolute error of the estimate with regard to computation cost for all methods. The ground truth value is obtained by running standard MC with $10^{8}$ samples. We can see that with more samples, the absolute errors decrease for all methods. Layering in time (MLMC) or in state (MFMC) alone results in error reduction compared to standard MC and subset simulation. By leveraging both, the proposed MLMF outperforms all other methods.